how do i solve the linear system:
x-2y=9
3x-y=17
I do not understand it because the y value does not cancel out, so what am i supposed to do with it?

Answer by jim_thompson5910(35256)

you can multiply both sides of equation 2 by -2 to get

3x-y=17
-2*(3x-y)=-2*17
-6x+2y = -34

so we now have this new system

x-2y=9
-6x+2y = -34


Add the equations

x-2y=9
-6x+2y = -34
------------
-5x+0y = -25

So -5x = -25 which means x = 5

Use this to find y.

x-2y=9
5-2y=9 ... plug in x = 5, solve for y
-2y = 9-5
-2y = 4
y = 4/(-2)
y = -2

So the ordered pair solution is (5,-2)